# Compositional Flows for 3D Molecule and Synthesis Pathway Co-design

### v1

@@ Abstract
Many generative applications, such as synthesis-based 3D molecular design, involve constructing compositional objects with continuous features. Here, we introduce Compositional Generative Flows (CGFlow), a novel framework that extends flow matching to generate objects in compositional steps while modeling continuous states. Our key insight is that modeling compositional state transitions can be formulated as a straightforward extension of the flow matching interpolation process. We further build upon the theoretical foundations of generative flow networks (GFlowNets), enabling reward-guided sampling of compositional structures. We apply CGFlow to synthesizable drug design by jointly designing the molecule’s synthetic pathway with its 3D binding pose. Our approach achieves state-of-the-art binding affinity on all 15 targets from the LIT-PCBA benchmark, and 5.8$\times$ improvement in sampling efficiency compared to 2D synthesis-based baseline. To our best knowledge, our method is also the first to achieve state of-art-performance in both Vina Dock (-9.38) and AiZynth success rate (62.2%) on the CrossDocked benchmark.
@@ 1 Introduction
In our experiments, we evaluate 3DSynthFlow on the task of designing synthesizable drugs directly within protein pockets. 3DSynthFlow achieves 5.8x sampling efficiency improvement, and state-of-the-art performance across all 15 targets in the LIT-PCBA benchmark (Tran-Nguyen et al., 2020) for binding affinity. 3DSynthFlow can be extended to pocket-conditional setting and achieve state of-art-performance in both Vina Dock (-9.38) and AiZynth success rate (62.2%) on the CrossDocked benchmark.
@@
  • We use our compositional framework to develop 3DSynthFlow for 3D molecule and synthesis pathway co-design. 3DSynthFlow achieves 5.8x improvement in sampling efficiency and demonstrate significant improvements in binding affinity and ligand efficiency on the LIT-PCBA benchmark.
@@ 3.3 Training objectives
We train CGFlow to jointly approximate the distribution of continuous states ${\mathcal{S}}$ and the sequential generation of the compositional structure ${\mathcal{C}}$ with two models: (1) a state flow model $p_{1|t}^{\theta}$ for updating continuous states, and (2) a compositional flow policy $\pi^{\theta}$ for sampling compositional components. Both models are conditioned on the entire object ${\bm{x}}_{t}=({\mathcal{C}}_{t},{\mathcal{S}}_{t})$ with self-conditioning $\hat{{\bm{x}}}_{1}^{t}=({\mathcal{C}}_{t},\hat{\mathcal{S}}_{1}^{t})$ at time $t$, where $\hat{\mathcal{S}}_{1}^{t}=(\hat{{\bm{S}}}^{(i)}_{1})_{i=1}^{k(t)}$ from the previous prediction (Chen et al., 2022). This formulation ensures their dependence on both the compositional structure and the continuous state of the object.
@@
Given a deterministic transition function and a fixed state flow model ODE, ${\bm{x}}_{t}=({\mathcal{C}}_{t},{\mathcal{S}}_{t})$ is uniquely determined by a given trajectory $\tau$ sampled from the compositional flow model $\pi^{\phi}$. This formulation views the interpolation with the state flow model $p_{1|t}^{\theta}$ as the transition of the state resulting from the sampled action by the compositional flow policy $\pi^{\phi}$111 Conceptually, it resembles an internal model of the world: used to predict the next state of the world as a function of the current state and action (Bryson & Ho, 1969). The state flow model weights are fixed to remain faithful to the data distribution. Then, TB loss is computed as:
@@
We provide the theoretical background for this derivation in Sec. B.
By enabling on-policy training, this approach generalizes beyond trajectories observed in the empirical data, as it can sample new data under the current policy. Alternatively, the compositional flow model can be trained using a maximum likelihood objective based on the data distribution (see Sec. A.3).
We refer readers to Sec. A.1 for a summary of the key steps in applying CGFlow to a new data domain. In the next section, we demostrate CGFlow’s application to 3D molecular generation and synthesis pathway design.
@@
State flow model predicts the docking pose of a molecule within a target protein pocket. Unlike typical docking, where the full 2D molecule is provided at $t=0$, the state flow model must predict the pose for the ligand that is sequentially constructed at discrete time intervals. Molecules for training the state flow model ${\bm{x}}$ are first decomposed into their compositional structure ${\mathcal{C}}$, which represents the sequentially composed synthesis pathway using Enamine reaction rules, and their continuous state ${\mathcal{S}}$, which represents the 3D coordinates refined over time. We discuss the motivation behind this formulation in Sec. E.1.2.
The state flow model is then trained using the objective described in Sec. 3.3.1 on the protein-ligand complex dataset, independent of the compositional flow model. The Semla architecture (Irwin et al., 2024), originally designed for 3D molecular generation, is used to parameterize the State flow model. We introduce new protein-encoding and protein and ligand message passing layers to enable protein-ligand modeling (See Sec. D.2.1).
Compositional flow model generates a synthesis pathway to construct a molecular structure. The action space for the composition flow model consists of all valid synthetic steps for a given compositional structure ${\mathcal{C}}_{t}$ (details in Sec. D.1). We modify the architecture from RxnFlow (Seo et al., 2024) to model the sampling policy for synthesis steps using 3D protein-ligand complexes as input (see Sec. D.2.2). The compositional flow model is trained online as described in Sec. 3.3.2.
@@ 5 Related Work
Diffusion for sequential data. Diffusion models have recently gained traction for generative modeling of sequential structures in diverse domains, spanning biological sequences (Campbell et al., 2024; Stark et al., 2024), videos (Ruhe et al., 2024), and language modeling (Lou et al., 2024). Campbell et al. (2023) propose a jump process for transitioning between different dimensional spaces to address the variable-dimension nature of data. To exploit the temporal causal dependency in sequences, Ruhe et al. (2024); Zhang et al. (2023) explore frame-level noise schedules for diffusion-based video generation for arbitrary-length frame rollout. Wu et al. (2023); Chen et al. (2024) apply similar ideas of training next-token prediction models while diffusing past ones for applications in planning and language modeling. Most similar to our work for molecule generation using diffusion models are methods that use a separate diffusion process for each sequentially added fragment (Peng et al., 2023; Ghorbani et al., 2023; Li et al., 2024).
We provide an extended related works for sequential diffusion for molecular generation and structure-based drug design in App. C.
Figure 4: Illustration of synthesis-based composition rules (a) reactant unit-based and (b) synthon unit-based (ours).
@@ 6 Experiments
Setup. We utilize 1.2M molecular fragments from the Enamine Catalog and 38 bimolecular Enamine synthesis protocols from Gao et al. (2024). To ensure tractability, molecular generation is limited to two synthesis steps, consistent with Enamine REAL Space guidelines (Grygorenko et al., 2020). To estimate synthesizability, we employ the retrosynthetic analysis tool AiZynthFinder (Genheden et al., 2020).
@@ 6.1 Pocket-specific optimization
Setup. We follow Seo et al. (2024) in both the reward function and evaluation protocol. To estimate the binding affinity, we use the GPU-accelerated docking tool UniDock (Yu et al., 2023). Each method generates up to 64,000 molecules for each protein target. To prevent reward hacking by increasing molecular size to improve the docking score, QED (Bickerton et al., 2012) is jointly optimized, and we set a generation of 40 heavy atoms. This ensure that the generated molecules have drug-like properties. Specifically, we optimize $R(x)=w_{1}\text{QED}(x)+w_{2}\widehat{\text{Vina}}(x)$, where $\widehat{\text{Vina}}(x)$ is a normalized docking score. The parameters $w_{1}$ and $w_{2}$ serve as conditions of multi-objective GFlowNets (Jain et al., 2023b), and are set to 0.5 for non-GFlowNet baselines.
@@
Figure 5: Number of discovered modes (satisfying Vina <-10 kcal/mol, QED >0.5, Sim <0.5) as a function of sampling budget for the first four LIT-PCBA targets for 3DSynthFlow (3D) vs RxnFlow (2D) across 2 seeds. Lower is better.
@@
Results. Table. 1 presents the Vina results for the first five targets, while the full property results for all targets are in Sec. F.8. 3DSynthFlow consistently outperforms all baselines across LIT-PCBA targets in affinity and ligand efficiency, demonstrating that co-designing 3D molecular structures alongside synthesis pathways enhances the discovery of high-affinity molecules. Furthermore, 3DSynthFlow achieves 5.8x sampling efficiency to 2D baselines, as detailed in Sec. 6.1.1. Molecules generated by 3DSynthFlow also exhibit a higher number of protein-ligand interactions, as quantified by PoseCheck, highlighting the benefits of 3D-aware modeling for target-based drug design (see Sec. F.5).
@@
6.1.1 3DSynthFlow improves sampling efficiency
We evaluate sample efficiency on the first 5 LIT-PCBA targets and report results in Fig. 5 and Table. 15. Diverse high-scoring modes were defined by QED >0.5, Vina <-10 kcal/mol [^3] Except for FEN1, where we use Vina threshold below -7 kcal/mol to maintain similar number of modes compared to the other four targets., and Tanimoto similarity <0.5 to any other mode. After sampling 10,000 molecules, 3DSynthFlow identified 5.8x number of diverse high-scoring modes compared to RxnFlow (316 vs 54.6). This enhanced sampling efficiency validates the effectiveness of 3DSynthFlow framework, and suggests a higher probability of experimental success.
@@
Table 2: Benchmark Results for Generative Methods. We report the average (Avg.) and median (Med.) values for each metric when available. Reference denotes known actives. For methods where only one value is available, the median is indicated as “–”.
@@
3D Reaction 3DSynthFlow (low $\beta$) 100.0% -9.00 -9.27 0.72 0.72 55.0% 54.5% 0.79 24
3DSynthFlow (med $\beta$) 100.0% -9.16 -9.41 0.73 0.74 56.6% 56.0% 0.76 24
3DSynthFlow (high $\beta$) 100.0% -9.38 -9.62 0.74 0.74 62.2% 63.0% 0.66 24
Setup. Our method generalizes to pocket-conditional generation problem setting (Peng et al., 2022; Guan et al., 2023; Schneuing et al., 2024a), enabling the design of binders for unseen pockets with a single model and no additional oracle calls. We follow the same pocket-conditional experimental setup used in TacoGFN (Shen et al., 2024) and RxnFlow (Seo et al., 2024). 3DSynthFlow adopt the pre-trained proxy from TacoGFN trained on CrossDock2020 train set, which leverages pharmacophore representation (Seo & Kim, 2024), to compute rewards for training a pocket-conditional policy.
@@
Baselines. We compare 3DSynthFlow with state-of-the-art distribution learning-based models trained on a synthesizable drug set, including the autoregressive model Pocket2Mol (Peng et al., 2022), and diffusion-based methods: TargetDiff (Guan et al., 2023), DiffSBDD (Schneuing et al., 2024b), DecompDiff (Guan et al., 2024), and MolCraft (Qu et al., 2024). We further include comparisons with optimization-based approaches: the fragment-based method TacoGFN (Shen et al., 2024) and the reaction-based method RxnFlow (Seo et al., 2024). To ensure a fair evaluation of distribution learning-based methods, we use the docking proxy trained on CrossDocked training set. To balance exploration and exploitation during sampling, we vary the reward exponentiation parameter $\beta$ of 3DSynthFlow: low $\beta$ (sampled from $\mathcal{U}(1,64)$), medium $\beta$ ($\mathcal{U}(32,64)$), and high $\beta$ ($\mathcal{U}(48,64)$).
Results. As shown in Table. 2, 3DSynthFlow achieves significant improvements in pocket-conditional generation, particularly in synthesizability and docking scores. In particular, 3DSynthFlow- high $\beta$ attains an average docking score of -9.38 kcal/mol, outperforming RxnFlow (-8.85) and state-of-the-art diffusion-based methods like MolCRAFT-large (-9.25) and DecompDiff (-8.35). We attribute this improvement largely to our explicit consideration of 3D co-design in the reaction-based generation framework.
6.2.1 3DSynthFlow achieves competitive synthesizability
3DSynthFlow also ensures high AiZynth success rate via synthesis-based generation, achieving success rates of 55.0%-62.2% across different $\beta$ values in the pocket-conditioned setting. This surpasses synthesis-based RxnFlow (34.8%), and TacoGFN (1.3%), which optimize for SA score.
The reported average AiZynth success rates of RxnFlow on LIT-PCBA is 65.4% and is significantly higher to their performance on CrossDocked, and consistently outperforms 3DSynthFlow (avg. 53.12%) (See Table. 14). It’s unclear where the inter-dataset variability arises from for RxnFlow. However, we note the synthesis success rate of our model is consistently strong for both CrossDock (55.0% - 62.2%) and LIT-PCBA benchmarks (53.12%).
SBDD methods such as MolCRAFT-large can attain strong binding affinity (-9.25) similar to 3DSynthFlow (-9.38), however their synthesis success rate is much lower (3.9% vs 62.2%). The main contribution of 3DSynthFlow is representing both the compositional nature for synthesis constraints and modeling 3D poses for binding. Both synthesizability and binding are both necessary requirement for experimental validation.

### v2

@@ Abstract
Many generative applications, such as synthesis-based 3D molecular design, involve constructing compositional objects with continuous features. Here, we introduce Compositional Generative Flows (CGFlow), a novel framework that extends flow matching to generate objects in compositional steps while modeling continuous states. Our key insight is that modeling compositional state transitions can be formulated as a straightforward extension of the flow matching interpolation process. We further build upon the theoretical foundations of generative flow networks (GFlowNets), enabling reward-guided sampling of compositional structures. We apply CGFlow to synthesizable drug design by jointly designing the molecule’s synthetic pathway with its 3D binding pose. Our approach achieves state-of-the-art binding affinity and synthesizability on all 15 targets from the LIT-PCBA benchmark, and 4.2$\times$ improvement in sampling efficiency compared to 2D synthesis-based baseline. To our best knowledge, our method is also the first to achieve state of-art-performance in both Vina Dock (-9.42) and AiZynth success rate (36.1%) on the CrossDocked2020 benchmark.
@@ 1 Introduction
In our experiments, we evaluate 3DSynthFlow on the task of designing synthesizable drugs directly within protein pockets. 3DSynthFlow achieves 4.2$\times$ sampling efficiency improvement, and state-of-the-art performance across all 15 targets in the LIT-PCBA benchmark (Tran-Nguyen et al., 2020) for binding affinity. 3DSynthFlow can be extended to pocket-conditional setting and achieve state of-art-performance in both Vina Dock (-9.42) and AiZynth success rate (36.1%) on the CrossDocked benchmark.
@@
  • We use our compositional framework to develop 3DSynthFlow for 3D molecule and synthesis pathway co-design. 3DSynthFlow achieves 4.2$\times$ improvement in sampling efficiency and demonstrate significant improvements in binding affinity and ligand efficiency on the LIT-PCBA benchmark.
@@ 3.3 Training objectives
We train CGFlow to jointly approximate the distribution of continuous states ${\mathcal{S}}$ and the sequential generation of the compositional structure ${\mathcal{C}}$ with two models: (1) a state flow model $p_{1|t}^{\theta}$ for updating continuous states, and (2) a compositional flow policy $\pi^{\theta}$ for sampling compositional components. Both models are conditioned on the object ${\bm{x}}_{t}=({\mathcal{C}}_{t},{\mathcal{S}}_{t})$ with self-conditioning $\hat{{\bm{x}}}_{1}^{t}=({\mathcal{C}}_{t},\hat{\mathcal{S}}_{1}^{t})$ at time $t$, where $\hat{\mathcal{S}}_{1}^{t}=(\hat{{\bm{S}}}^{(i)}_{1})_{i=1}^{k(t)}$ from the previous prediction (Chen et al., 2022). This ensures their dependence on both the compositional structure and the continuous state of the object.
@@
Given a deterministic transition function and a fixed state flow model ODE, ${\bm{x}}_{t}=({\mathcal{C}}_{t},{\mathcal{S}}_{t})$ is uniquely determined by a given trajectory $\tau$ sampled from the compositional flow model $\pi^{\phi}$. This formulation views the interpolation with the state flow model $p_{1|t}^{\theta}$ as the transition of the state resulting from the sampled action by the compositional flow policy $\pi^{\phi}$ [^1] Conceptually, it resembles an internal model of the world: used to predict the next state of the world as a function of the current state and action (Bryson & Ho, 1969). The state flow model weights are fixed to remain faithful to the data distribution.. Then, TB loss is computed as:
@@
We provide the theoretical background in Sec. B.
By enabling online training, this approach generalizes beyond the empirical data, as it can sample new data under the current policy. Alternatively, the compositional flow model can be trained using a maximum likelihood objective based on the data distribution (see Sec. A.3).
We refer readers to Sec. A.1 for a summary of the key steps in applying CGFlow to a new data domain. In the next section, we demonstrate CGFlow’s application to 3D molecular generation and synthesis pathway design.
Figure 4: Illustration of synthesis-based composition rules. (a) reactant unit-based and (b) synthon unit-based (ours).
@@
State flow model predicts the docking pose of a molecule within a target protein pocket. Unlike typical docking, where the full 2D molecule is provided at $t=0$, the state flow model must predict the pose for the ligand that is sequentially constructed at discrete time intervals. Molecules for training the state flow model ${\bm{x}}$ are first decomposed into their compositional structure ${\mathcal{C}}$, which represents the sequentially composed synthesis pathway using Enamine reaction rules, and their continuous state ${\mathcal{S}}$, which represents the 3D coordinates refined over time. To ensure atoms not present in the molecule do not appear during decomposition (i.e., leaving group for reactions), we adopt a synthon-based representation rather than a reactant-based representation as Fig. 4. We discuss the motivation behind this formulation in Sec. E.1.2.
The state flow model is then trained using the objective described in Sec. 3.3.1 on the protein-ligand complex dataset, independent of the compositional flow model. The Semla architecture (Irwin et al., 2024), originally designed for 3D molecular generation, is used to parameterize the State flow model. We introduce new hierarchical Semla for memory-efficient protein encoding, and protein-ligand message passing to enable protein-ligand modeling (See Sec. D.2.1).
Compositional flow model generates a synthesis pathway to construct a molecular structure. The action space for the composition flow model consists of all valid synthetic steps for a given structure ${\mathcal{C}}_{t}$ (see Sec. D.1). We modify the architecture from RxnFlow (Seo et al., 2024) to model the sampling policy for synthesis steps using 3D protein-ligand complexes as input (see Sec. D.2.2). The compositional flow model is trained online as described in Sec. 3.3.2.
@@ 5 Related Work
Diffusion for sequential data. Diffusion models have recently gained traction for generative modeling of sequential structures in diverse domains, spanning biological sequences (Campbell et al., 2024; Stark et al., 2024), videos (Ruhe et al., 2024), and language modeling (Lou et al., 2024). Campbell et al. (2023) propose a jump process for transitioning between different dimensional spaces to address the variable-dimension nature of data. To exploit the temporal causal dependency in sequences, Ruhe et al. (2024); Zhang et al. (2023) explore frame-level noise schedules for diffusion-based video generation for arbitrary-length frame rollout. Wu et al. (2023); Chen et al. (2024) apply similar ideas of training next-token prediction models while diffusing past ones for applications in planning and language modeling. Most similar to our work for molecule design using diffusion models are methods that use a separate diffusion process for each sequentially added fragment (Peng et al., 2023; Ghorbani et al., 2023; Li et al., 2024).
We provide an extended related works for sequential diffusion for molecular generation and structure-based drug design in Sec. C.
@@ 6 Experiments
Setup. To construct synthetic pathway, we utilize 38 bimolecular Enamine synthesis protocols from Gao et al. (2024). We use Enamine’s 1.2M-size catalog set and 300k-size stock set in Sec. 6.1 and Sec. 6.2, respectively. For stock set, we filtered out molecules with drug-likeness score lower than 60 (Lee et al., 2022). To ensure tractability, molecular generation is limited to two synthesis steps, consistent with Enamine REAL Space (Grygorenko et al., 2020). To estimate synthesizability, we employ the retrosynthetic analysis tool AiZynthFinder (Genheden et al., 2020).
@@ 6.1 Pocket-specific optimization
Setup. We follow Seo et al. (2024) in both the reward function and evaluation protocol. To estimate the binding affinity, we use the GPU-accelerated docking tool UniDock (Yu et al., 2023). Each method generates up to 64,000 molecules for each protein target. To prevent reward hacking by increasing molecular size to improve the docking score, QED (Bickerton et al., 2012) is jointly optimized, and we set a generation of 40 heavy atoms. The reward function for multi-objective optimization is described in Sec. E.2.
@@
Figure 5: Number of discovered modes (satisfying Vina <-10 kcal/mol, QED >0.5, Sim <0.5) as a function of sampling budget for the first four LIT-PCBA targets for 3DSynthFlow (3D) vs RxnFlow (2D) across 3 seeds. Higher is better.
@@
Results. Table. 1 presents the Vina results for the first five targets, while the full property results for all targets are in Sec. F.7. 3DSynthFlow consistently outperforms all baselines across LIT-PCBA targets in affinity and ligand efficiency, demonstrating that co-designing 3D molecular structures alongside synthesis pathways enhances the discovery of high-affinity molecules. Table. 3 shows molecules generated by 3DSynthFlow also exhibit a higher number of protein-ligand interactions, as quantified by PoseCheck, highlighting the benefits of 3D-aware modeling for target-aware drug design (see full results in Sec. F.5).
Furthermore, Fig. 5 and Table. 16 presents the sampling efficiency for the first 5 LIT-PCBA targets. Diverse high-scoring modes were defined by QED>0.5, Vina<-10 kcal/mol [^3] Except for FEN1, where we use Vina below -7 kcal/mol to maintain similar number of modes compared to the other targets., and Tanimoto similarity<0.5 to any other mode. After sampling 64,000 molecules, 3DSynthFlow identified 4.2x number of diverse high-scoring modes compared to RxnFlow (4432.3 vs 1062.6). This enhanced sampling efficiency validates the effectiveness of 3DSynthFlow framework, and suggests a higher probability of experimental success.
Method Success Rate (%, $\uparrow$) Synthesis Steps ($\downarrow$)
FragGFN+SA 03.52 3.74
SynNet 47.50 3.45
BBAR 17.92 3.68
SynFlowNeta 54.60 2.55
SynFlowNetb 58.38 2.47
RGFN 47.43 2.46
RxnFlow 65.35 2.17
3DSynthFlow 68.58 2.39
Table 3: PoseCheck Averages and standard deviations over 4 runs, for the top 100 diverse modes for the first 5 LIT-PCBA pocket. The results for all protein-ligand interactions are reported in Appendix. F.5. The best results in each column are in bold.
Method H-Bond Acceptors H-Bond Donors Sum
SynFlowNeta 0.22 ($\pm$ 0.01) 0.11 ($\pm$ 0.01) 0.33
SynFlowNetb 0.22 ($\pm$ 0.03) 0.10 ($\pm$ 0.01) 0.32
RGFN 0.19 ($\pm$ 0.01) 0.11 ($\pm$ 0.01) 0.30
RxnFlow 0.22 ($\pm$ 0.00) 0.10 ($\pm$ 0.01) 0.32
3DSynthFlow 0.33 ($\pm$ 0.06) 0.17 ($\pm$ 0.03) 0.50
Table. 2 shows that 3DSynthFlow attains comparable synthesizability compared to its 2D baselines, demonstrating its generated molecules is likely to be synthesizable. This ensures that the high-affinity compounds identified by our model are not merely theoretical but represent viable candidates for experimental validation and further development.
@@
Table 4: Benchmark Results for Generative Methods. We report the average (Avg.) and median (Med.) values for each metric when available. Time indicates the average duration to generate 100 molecules. Reference denotes known actives. For methods where only one value is available, the median is indicated as “–”.
@@
3D Reaction 3DSynthFlow (low $\beta$) 099.9% -9.14 -9.38 0.69 0.69 36.2% 37.0% 0.78 6
3DSynthFlow (med $\beta$) 099.9% -9.30 -9.62 0.72 0.71 35.1% 36.0% 0.74 6
3DSynthFlow (high $\beta$) 100.0% -9.42 -9.61 0.73 0.73 36.1% 36.0% 0.69 6
Setup. Our method generalizes to pocket-conditional generation problem setting (Peng et al., 2022; Guan et al., 2023; Schneuing et al., 2024a), enabling the design of binders for unseen pockets with a single model and no additional oracle calls. We follow the same pocket-conditional experimental setup and reward function used in TacoGFN (Shen et al., 2024) and RxnFlow (Seo et al., 2024) (see Sec. E.2. 3DSynthFlow adopt the pre-trained proxy from TacoGFN trained on CrossDock2020 train set, which leverages pharmacophore representation (Seo & Kim, 2024), to compute rewards for training a pocket-conditional policy.
@@
Baselines. We compare 3DSynthFlow with state-of-the-art distribution learning-based models trained on a synthesizable drug set, including the autoregressive model Pocket2Mol (Peng et al., 2022), and diffusion-based methods: TargetDiff (Guan et al., 2023), DiffSBDD (Schneuing et al., 2024b), DecompDiff (Guan et al., 2024), and MolCraft (Qu et al., 2024). We further include comparisons with optimization-based approaches: the fragment-based method TacoGFN (Shen et al., 2024) and the reaction-based method RxnFlow (Seo et al., 2024). To ensure a fair evaluation of distribution learning-based methods, we use the docking proxy trained on CrossDocked training set. To balance exploration and exploitation during sampling, we vary the reward exponentiation parameter $\beta$ of 3DSynthFlow: low $\beta$ (sampled from $\mathcal{U}(1,64)$), medium $\beta$ ($\mathcal{U}(16,64)$), and high $\beta$ ($\mathcal{U}(32,64)$).
Results. As shown in Table. 4, 3DSynthFlow achieves improvements in pocket-conditional generation in synthesizability and docking scores. In particular, 3DSynthFlow- high $\beta$ attains an average docking score of -9.42 kcal/mol, outperforming RxnFlow (-8.85) and state-of-the-art diffusion-based methods like MolCRAFT-large (-9.25) and DecompDiff (-8.35). We attribute this improvement largely to our explicit consideration of 3D co-design in the reaction-based generation framework. However, due to the limited accuracy of the pre-trained proxy model for docking score reward, we find the improvement in docking score less significant than the pocket-specific setting.
3DSynthFlow surpasses the 2D reaction-based GFlowNets, TacoGFN and RxnFlow, while maintaining a highly comparable computational cost. We attribute this to a fundamental design difference: while the 2D baselines are conditioned on 1D sequence-based representations, 3DSynthFlow leverages the full 3D geometry of the binding site. This result indicates that 3D interaction modeling is essential, particularly for achieving robust performance across diverse targets.
SBDD methods such as MolCRAFT-large can attain strong binding affinity (-9.25) similar to 3DSynthFlow (-9.42), however their synthesis success rate is much lower (3.9% vs 36.1%). The main contribution of 3DSynthFlow is representing both the compositional nature for synthesis constraints and modeling 3D poses for binding.